# Weak-Coupling Limit of the Lattice Nonlinear Schrödinger Integral Equation

### 10 Mar 2026 v1

@@ Abstract
We study the ground-state integral equation of the quantum lattice nonlinear Schrödinger model—equivalently the isotropic Heisenberg XXX spin chain with spin s=1s=-1—in the weak-coupling limit. Unlike the continuous Lieb–Liniger equation, whose driving term is a constant, the lattice equation is doubly singular: both the driving term and the integral kernel degenerate into $\delta$-functions as $\kappa\to 0$. We develop a matched asymptotic expansion with three regions—inner, outer, and edge. The Fourier transform of the rescaled inner solution is exactly the Bose–Einstein distribution, and the peak density diverges logarithmically with a constant $C$, which we determine analytically via two independent routes and confirm numerically. An exact duality with the Love integral equation for the circular disc capacitor yields the total density expansion. We prove an identity for the inner energy, allowing us to obtain the ground-state energy per site. From the Wiener–Hopf factorisation of the edge boundary layer, we identify the instanton action and predict a resurgent transseries structure.
@@ 1 Introduction
The discovery that the Lieb–Liniger expansion is non-Borel summable [65] sparked a programme of resurgent analysis in integrable quantum field theories. The key tool is the Wiener–Hopf factorisation of the integral equation’s symbol (developed in this context in [88, 89]), which gives systematic access to both perturbative and non-perturbative sectors. Extending such an analysis to renormalons in integrable field theories has also been the subject of investigation [67, 66], not to mention the Hubbard model [68], and the energy gap problem [69]. Important developments were carried out on the $1/N$ expansion [73] and detailed analyses of the $O(4)$ and $O(3)$ sigma models [1, 5, 3]. More recently, Bajnok, Balog, and Vona [6] obtained the complete transseries for conserved charges. These developments are reviewed in [2, 19].
The quantum inverse scattering method (QISM) [81, 82] provides the algebraic framework for constructing integrable quantum models on the lattice (see also [23, 24, 25, 26] for reviews). The quantum nonlinear Schrödinger (NLS) model on a lattice was constructed within this framework [39, 40, 52, 82], with local Hamiltonians worked out in [85]. The lattice NLS provides an integrable discretisation of the continuous NLS field theory, preserving the full integrability structure at the quantum level [49, 12]. The model is closely related to the $q$-boson hopping model [11] and it was diagonalised in [87], while Dorlas [20] established completeness of the Bethe ansatz eigenstates for the NLS model.
@@
The same integral equation (2) appears in high-energy quantum chromodynamics: the $\kappa\to 0$ limit of the XXXs=-1 chain corresponds to the weak-coupling/high-density regime of reggeized gluon dynamics. Furthermore, it was discovered [60, 61, 62, 63] that the BFKL pomeron equation—governing small-$x$ parton evolution [27, 53, 54, 7]—possesses a hidden integrable structure equivalent to a non-compact Heisenberg spin chain. Faddeev and Korchemsky [22] proved the complete integrability of the BFKL Hamiltonian and identified it with the XXXs=0 spin chain, while Korchemsky [45, 46, 47] developed the Bethe ansatz for the QCD pomeron and studied its quasiclassical limit (see also [8, 48] for reviews). Finally, de Vega and Lipatov [15] extended the analysis to multi-reggeon interactions. In this context, the logarithmic scaling we find is qualitatively consistent with the logarithmic energy dependence characteristic of BFKL dynamics, though a precise quantitative comparison requires further work (see Section 2).
@@
In this paper we develop a systematic matched asymptotic expansion for equation (2) as $\kappa\to 0$, exploiting three distinct regions of the rescaled spectral variable $\xi=\lambda/\kappa$. The methods draw on classical techniques of singular perturbation theory [21, 9, 36, 55] and on the Wiener–Hopf factorisation technique [72, 50, 33] as applied to convolution integral equations on finite intervals [80]. The matched asymptotic expansion is formal in the sense of classical singular perturbation theory; the leading-order results are confirmed analytically by the Wiener–Hopf derivation of Section 6.2 and verified numerically throughout. Our main findings are: i) in the inner region ($|\xi|\lesssim 1$), the equation reduces to a Lieb–Liniger-type equation on the expanding domain [Q,Q][-Q,Q] with $Q=q/\kappa\to\infty$. On the full line, the Fourier transform of the rescaled solution is exactly the Bose–Einstein distribution $\hat{\tilde{\rho}}(p)=1/(e^{|p|}-1)$—a function familiar from the Yang–Yang thermodynamic Bethe ansatz [90, 92]—whose infrared singularity produces a logarithmic divergence at the origin. We determine the constant analytically as $C=(\gamma_{E}+\log 2)/\pi$ via two independent routes: a mode-counting argument combined with a digamma representation of the solution, and a direct Wiener–Hopf derivation using the spectral response function of the truncated operator. Both are confirmed numerically by Richardson extrapolation; ii) in the outer region ($1\ll|\xi|\ll Q$), the solution reduces to a uniform Fermi sea with $\tilde{\rho}_{\mathrm{bulk}}=1/2$, which dominates the total density. Using a duality with the Love integral equation [64, 28], together with the self-consistency identity $f(0)=1+D(Q)$ and a novel integral identity for the digamma function, we derive the total density expansion $D(Q)=Q+(2\pi)^{-1}\log Q+b+\cdots\,$; iii) In the edge boundary layer of width $\mathcal{O}(1)$ in the rescaled variable near $\xi=\pm Q$, the equation reduces to a half-line Wiener–Hopf problem with symbol $\Sigma=1-e^{-|p|}$. We obtain the factorisation of this symbol in terms of gamma functions and identify the Wiener–Hopf constant $A_{\mathrm{WH}}=2$. The structure of the edge problem is closely related to the Kirchhoff–Maxwell analysis of fringing fields in the circular disc capacitor [43, 37, 28, 4]; iv) We prove the identity $E_{\mathrm{inner}}(Q)=2\pi\tilde{\rho}(0;Q)-2$, valid for all $Q>0$, which reduces the energy calculation to the peak density. The physical ground-state energy per site is then $e(\kappa)=-2[\log(2q/\kappa)+\gamma_{E}-1]/\kappa+\cdots\,$, scaling as $e(\kappa)\sim-2\log(1/\kappa)/\kappa$—qualitatively different from the Lieb–Liniger energy $e(\gamma)\sim\gamma$ at weak coupling. All analytical predictions are confirmed by numerical computations.
@@ 2.1 Lattice nonlinear Schrödinger model
The same non-compact $\mathfrak{sl}(2)$ representations appear in the integrable structures of high-energy quantum chromodynamics [60, 61, 62, 63].  [22] proved that the BFKL pomeron Hamiltonian—governing the small-$x$ evolution of parton distributions [27, 53, 54, 7]—is equivalent to the Hamiltonian of a non-compact Heisenberg spin chain with $\mathfrak{sl}(2,\mathbb{R})$ symmetry, and [45, 46, 47] developed the Bethe ansatz for this QCD spin chain (see also [8, 48] for reviews). Our integral equation (2) governs the ground state in the high-density regime, and the logarithmic scaling $\tilde{\rho}(0;Q)\sim(\log Q)/\pi$ that we find is qualitatively consistent with the $\log(1/x)$ dependence characteristic of BFKL dynamics. However, since the BFKL pomeron corresponds to spin $s=0$ in the Faddeev–Korchemsky identification [22] whereas our model has s=1s=-1, a precise quantitative mapping requires a careful continuation in the spin parameter, which is left for future work.
@@ 2.2 Bethe ansatz and thermodynamic limit
where the Lorentzian kernel is $K(\lambda;\,\kappa)=2\kappa/(\kappa^{2}+\lambda^{2})$, normalised so that $\int_{-\infty}^{\infty}K(\lambda;\,\kappa)\,d\lambda=2\pi$ for all $\kappa>0$. This is a Fredholm integral equation of the second kind on the symmetric interval [q,q][-q,q], with a Lorentzian kernel that is symmetric, positive, and of convolution type.
@@
The kernel appearing in the energy integral is the same function $K(\lambda;\,\kappa)$ that serves as both the driving term and the scattering kernel in (19)—a coincidence specific to the lattice NLS model that will have important consequences [^2] See the identity $E_{\mathrm{inner}}=2\pi\tilde{\rho}(0)-2$ in Section 7. The total momentum per site is $P=0$ for the ground state, by parity. Excited states are described by particle–hole excitations above the Fermi sea, with dressed energy and dressed momentum satisfying linear integral equations of the same type as (19), with different driving terms [49, 90]. The excitation spectrum and thermodynamics of the XXXs=-1 chain were recently studied [93]. In this manuscript, we restrict ourselves to the ground-state.
@@ 3.1 Rescaled inner equation
which “zoom in” on the region $|\lambda|\sim\kappa$ where the driving term is concentrated. Note that the rescaling preserves the particle density:
@@
since $\tilde{\rho}(\xi)\,d\xi=\kappa\rho(\kappa\xi)\cdot d\lambda/\kappa=\rho(\lambda)\,d\lambda$. Under this change of variables, equation (19) becomes
@@
This is a Lieb–Liniger-type equation on the expanding domain [Q,Q][-Q,Q], but with the constant driving term $1$ replaced by the Lorentzian
@@
and the coupling parameter set to $\kappa=1$. The limit $\kappa\to 0$ at fixed $q$ corresponds to $Q=q/\kappa\to\infty$.
The physical observables transform as:
@@ 3.2 Fourier analysis and logarithmic divergence
Figure 1: Numerical solution of the rescaled integral equation (25) for $Q=20$, $50$, $100$, and $200$. (a) Rescaled density $\tilde{\rho}(\xi;\,Q)$ plotted against $\xi/Q$, showing the full domain [Q,Q][-Q,Q]. All curves share the same outer-region (Fermi-sea) plateau at $\tilde{\rho}_{\mathrm{bulk}}=1/2$ (dashed line), while the central Bose–Einstein peak grows logarithmically with $Q$. The edge boundary layers at $\xi/Q=\pm 1$ are visible as the rapid drop from the plateau to zero. (b) Zoom into the inner region $|\xi|\lesssim\mathcal{O}(1)$, plotted against the unscaled variable $\xi$. The peak height at $\xi=0$ increases as $\tilde{\rho}(0;\,Q)\sim(\log Q)/\pi+C$ (see (42)), with the logarithmic growth clearly visible across the four values of $Q$. The cusp-like shape reflects the $1/|p|$ infrared singularity of the Bose–Einstein distribution (34) in Fourier space, and the universal profile is controlled by the digamma function $\operatorname{Re}\psi(1+i\xi)$. In panel (b), the dashed curves show the analytical inner approximation $\tilde{\rho}(\xi;Q)\approx[\log(2Q)-\operatorname{Re}\psi(1+i\xi)]/\pi$ from (45), which matches the numerical solution to graphical accuracy for $|\xi|\lesssim Q/2$.
@@ 3.3 Eigenvalue analysis of the truncated kernel
where $C$ is a constant determined by the detailed structure of the finite-interval problem, including the boundary layers near $\xi=\pm Q$. Further analytical details on the spectral properties of $\mathcal{K}_{Q}$ are given in Appendix A; the closing of the spectral gap is illustrated in Figure 2. The integral equation (25) is also closely related to the classical Love integral equation [64] for the circular disc capacitor; this connection, and the resulting duality that links the lattice NLS density to the Love solution, is developed in detail in Section 4.2.
@@ 4.1 Outer solution and Fermi sea
Since $\int_{-\infty}^{\infty}K(x)\,dx=2\pi$ and the solution varies slowly on the $\mathcal{O}(1)$ scale of the kernel, the convolution integral evaluates to $\approx 2\pi\tilde{\rho}$, and the equation is trivially satisfied for any constant $\tilde{\rho}$.
@@
Numerically, sampling $\tilde{\rho}$ in the outer region ($|\xi|\in[5,Q-5]$) and computing $D(Q)/(2Q)$ both converge to $1/2$ as $Q\to\infty$.
@@ 4.2 Duality with the Love integral equation
The integral equation (25) is closely related to the Love integral equation [64] for the electrostatic potential of two coaxial circular conducting discs separated by a distance $2/Q$ (in appropriate units). The Love equation reads
@@
$(I-\mathcal{L}_{Q})\tilde{\rho}=g/(2\pi),\qquad g(\xi)=2/(1+\xi^{2})\,$, (52)
while the Love equation is $(I-\mathcal{L}_{Q})f=1$. The two equations share the same operator $I-\mathcal{L}_{Q}$ but have different right-hand sides: the Lorentzian $g/(2\pi)$ in the lattice NLS case, versus the constant function $\mathbf{1}$ in the Love case.
@@
$D(Q)=\langle\mathbf{1},\tilde{\rho}\rangle=\bigl\langle(I-\mathcal{L}_{Q})^{-1}\mathbf{1},\,g/(2\pi)\bigr\rangle=\left\langle f,\,\tfrac{g}{2\pi}\right\rangle=\frac{1}{\pi}\int_{-Q}^{Q}\frac{f(\xi)}{1+\xi^{2}}\,d\xi\,$, (53)
@@
Numerical verification: solving both equations independently at $Q=100$ with $N=1400$-point Gauss–Legendre quadrature, the two sides of (53) agree to $2.7\times 10^{-13}$, consistent with double-precision roundoff error propagated through the $N\times N$ linear solve.

### 22 Mar 2026 v2

@@ Abstract
We study the ground-state integral equation of the quantum lattice nonlinear Schrödinger model—equivalently the isotropic Heisenberg XXX spin chain with spin s=1s=-1—in the weak-coupling limit. Unlike the continuous Lieb–Liniger equation, whose driving term is a constant, the lattice equation is doubly singular: both the driving term and the integral kernel degenerate into $\delta$-functions as $\kappa\to 0$. We develop a matched asymptotic expansion with three regions—inner, outer, and edge. We show that the Fourier transform of the rescaled inner solution is exactly the Bose–Einstein distribution, and the peak density diverges logarithmically with a constant $C$, which we determine analytically via two independent routes and confirm numerically. A duality with the Love integral equation for the circular disc capacitor yields the total density expansion. We prove an identity for the inner energy, allowing us to obtain the ground-state energy per site. From the Wiener–Hopf factorisation of the edge boundary layer, we identify the instanton action and predict a resurgent transseries structure.
@@ 1 Introduction
The discovery that the Lieb–Liniger expansion is non-Borel summable [65] sparked a programme of resurgent analysis in integrable quantum field theories. The key tool is the Wiener–Hopf factorisation of the integral equation’s symbol (developed in this context in [88, 89]), which gives systematic access to both perturbative and non-perturbative sectors. Extending such an analysis to renormalons in integrable field theories has also been the subject of investigation [67, 66], not to mention the Hubbard model [68], and the energy gap problem [69]. Important developments were carried out on the $1/N$ expansion [73] and detailed analyses of the $O(4)$ and $O(3)$ sigma models [1, 5, 3]. More recently, [6] obtained the complete transseries for conserved charges. These developments are reviewed in [2, 19].
The quantum inverse scattering method (QISM) [81, 82] provides the algebraic framework for constructing integrable quantum models on the lattice (see also [23, 24, 25, 26] for reviews). The quantum nonlinear Schrödinger (NLS) model on a lattice was constructed within this framework [39, 40, 52, 82], with local Hamiltonians worked out in [85]. The lattice NLS provides an integrable discretisation of the continuous NLS field theory, preserving the full integrability structure at the quantum level [49, 12]. The model is closely related to the $q$-boson hopping model [11] and it was diagonalised in [87], while [20] established completeness of the Bethe ansatz eigenstates for the NLS model.
@@
The same integral equation (2) appears in high-energy quantum chromodynamics: the $\kappa\to 0$ limit of the XXXs=-1 chain corresponds to the weak-coupling/high-density regime of reggeized gluon dynamics. Furthermore, it was discovered [60, 61, 62, 63] that the BFKL pomeron equation—governing small-$x$ parton evolution [27, 53, 54, 7]—possesses a hidden integrable structure equivalent to a non-compact Heisenberg spin chain. Faddeev and Korchemsky [22] proved the complete integrability of the BFKL Hamiltonian and identified it with the XXXs=0 spin chain, while Korchemsky [45, 46, 47] developed the Bethe ansatz for the QCD pomeron and studied its quasiclassical limit (see also [8, 48] for reviews). Finally, [15] extended the analysis to multi-reggeon interactions. In this context, the logarithmic scaling we find is qualitatively consistent with the logarithmic energy dependence characteristic of BFKL dynamics, though a precise quantitative comparison requires further work (see Section 2).
@@
In this paper we develop a systematic matched asymptotic expansion for equation (2) as $\kappa\to 0$, exploiting three distinct regions of the rescaled spectral variable $\xi=\lambda/\kappa$. The methods draw on classical techniques of singular perturbation theory [21, 9, 36, 55] and on the Wiener–Hopf factorisation technique [72, 50, 33] as applied to convolution integral equations on finite intervals [80]. The matched asymptotic expansion is formal in the sense of classical singular perturbation theory; the leading-order results are confirmed analytically by the Wiener–Hopf derivation of Section 6.2 and verified numerically throughout. Our main findings are: i) in the inner region ($|\xi|\lesssim 1$), the equation reduces to a Lieb–Liniger-type equation on the expanding domain [Q,Q][-Q,Q] with $Q=q/\kappa\to\infty$. On the full line, the Fourier transform of the rescaled solution is exactly the Bose–Einstein distribution $\hat{\tilde{\rho}}(p)=1/(e^{|p|}-1)$—a function familiar from the Yang–Yang thermodynamic Bethe ansatz [90, 92]—whose infrared singularity produces a logarithmic divergence at the origin. We determine the constant analytically as $C=(\gamma_{E}+\log 2)/\pi$ via two independent routes: a mode-counting argument combined with a digamma representation of the solution, and a direct Wiener–Hopf derivation using the spectral response function of the truncated operator. Both are confirmed numerically; ii) in the outer region ($1\ll|\xi|\ll Q$), the solution reduces to a uniform Fermi sea with $\tilde{\rho}_{\mathrm{bulk}}=1/2$, which dominates the total density. Using a duality with the Love integral equation [64, 28], together with the self-consistency identity $f(0)=1+D(Q)$ and a novel integral identity for the digamma function, we derive the total density expansion $D(Q)=Q+(2\pi)^{-1}\log Q+b+\cdots\,$; iii) In the edge boundary layer of width $\mathcal{O}(1)$ in the rescaled variable near $\xi=\pm Q$, the equation reduces to a half-line Wiener–Hopf problem with symbol $\Sigma=1-e^{-|p|}$. We obtain the factorisation of this symbol in terms of gamma functions and identify the Wiener–Hopf constant $A_{\mathrm{WH}}=2$. The structure of the edge problem is closely related to the Kirchhoff–Maxwell analysis of fringing fields in the circular disc capacitor [43, 37, 28, 4]; iv) We prove the identity $E_{\mathrm{inner}}(Q)=2\pi\tilde{\rho}(0;Q)-2$, valid for all $Q>0$, which reduces the energy calculation to the peak density. The physical ground-state energy per site is then $e(\kappa)=-2[\log(2q/\kappa)+\gamma_{E}-1]/\kappa+\cdots\,$, scaling as $e(\kappa)\sim-2\log(1/\kappa)/\kappa$—qualitatively different from the Lieb–Liniger energy $e(\gamma)\sim\gamma$ at weak coupling. All analytical predictions are confirmed by numerical computations.
@@ 2.1 Lattice nonlinear Schrödinger model
The same non-compact $\mathfrak{sl}(2)$ representations appear in the integrable structures of high-energy QCD [60, 61, 62, 63].  [22] proved that the BFKL pomeron Hamiltonian—governing the small-$x$ evolution of parton distributions [27, 53, 54, 7]—is equivalent to the Hamiltonian of a non-compact Heisenberg spin chain with $\mathfrak{sl}(2,\mathbb{R})$ symmetry, and [45, 46, 47] developed the Bethe ansatz for this QCD spin chain (see also [8, 48] for reviews). Our integral equation (2) governs the ground state in the high-density regime, and the logarithmic scaling $\tilde{\rho}(0;Q)\sim(\log Q)/\pi$ that we find is qualitatively consistent with the $\log(1/x)$ dependence characteristic of BFKL dynamics. However, since the BFKL pomeron corresponds to spin $s=0$ in the Faddeev–Korchemsky identification [22] whereas our model has s=1s=-1, a precise quantitative mapping requires a careful continuation in the spin parameter, which is left for future work.
@@ 2.2 Bethe ansatz and thermodynamic limit
where the Lorentzian kernel is $K(\lambda;\,\kappa)=2\kappa/(\kappa^{2}+\lambda^{2})$, normalised so that
$\int_{-\infty}^{\infty}K(\lambda;\,\kappa)\,d\lambda=2\pi$ (20)
for all $\kappa>0$. This is a Fredholm integral equation of the second kind on the symmetric interval [q,q][-q,q], with a Lorentzian kernel that is symmetric, positive, and of convolution type.
@@
The kernel appearing in the energy integral is the same function $K(\lambda;\,\kappa)$ that serves as both the driving term and the scattering kernel in (19)—a coincidence specific to the lattice NLS model that will have important consequences [^2] See the identity (124) in Section 7.. The total momentum per site is $P=0$ for the ground state, by parity. Excited states are described by particle–hole excitations above the Fermi sea, with dressed energy and dressed momentum satisfying linear integral equations of the same type as (19), with different driving terms [49, 90]. The excitation spectrum and thermodynamics of the XXXs=-1 chain were recently studied [93]. In this manuscript, we restrict ourselves to the ground-state.
@@ 3.1 Rescaled inner equation
which zoom in on the region $|\lambda|\sim\kappa$ where the driving term is concentrated. Note that the rescaling preserves the particle density
@@
since
$\tilde{\rho}(\xi)\,d\xi=\kappa\rho(\kappa\xi)\cdot\frac{d\lambda}{\kappa}=\rho(\lambda)\,d\lambda\,.$ (26)
Under this change of variables, equation (19) becomes
@@
This is a Lieb–Liniger-type equation on the expanding domain [Q,Q][-Q,Q], but with the constant driving term replaced by the Lorentzian
@@
and the coupling parameter set to $\kappa=1$. The limit $\kappa\to 0$ at fixed $q$ corresponds to $Q=q/\kappa\to\infty$. The physical observables transform as:
@@ 3.2 Fourier analysis and logarithmic divergence
Figure 1: Numerical solution of the rescaled integral equation (27) for $Q=20$, $50$, $100$, and $200$. (a) Rescaled density $\tilde{\rho}(\xi;\,Q)$ plotted against $\xi/Q$, showing the full domain [Q,Q][-Q,Q]. All curves share the same outer-region (Fermi-sea) plateau at $\tilde{\rho}_{\mathrm{bulk}}=1/2$ (dashed line), while the central Bose–Einstein peak grows logarithmically with $Q$. The edge boundary layers at $\xi/Q=\pm 1$ are visible as the rapid drop from the plateau to zero. (b) Zoom into the inner region $|\xi|\lesssim\mathcal{O}(1)$, plotted against the rescaled variable $\xi$. The peak height at $\xi=0$ increases as $\tilde{\rho}(0;\,Q)\sim(\log Q)/\pi+C$ (see (44)), with the logarithmic growth clearly visible across the four values of $Q$. The cusp-like shape reflects the $1/|p|$ infrared singularity of the Bose–Einstein distribution (36) in Fourier space, and the universal profile is controlled by the digamma function $\operatorname{Re}\psi(1+i\xi)$. In panel (b), the dashed curves show the analytical inner approximation $\tilde{\rho}(\xi;Q)\approx[\log(2Q)-\operatorname{Re}\psi(1+i\xi)]/\pi$ from (47), which matches the numerical solution to graphical accuracy for $|\xi|\lesssim Q/2$.
@@ 3.3 Eigenvalue analysis of the truncated kernel
where $C$ is a constant determined by the detailed structure of the finite-interval problem, including the boundary layers near $\xi=\pm Q$. Further analytical details on the spectral properties of $\mathcal{K}_{Q}$ are given in Appendix A; the closing of the spectral gap is illustrated in Figure 2. Numerical computation (Figure 2b) shows that $Q\cdot\Delta_{0}$ is not constant but grows slowly, well fitted by $Q\cdot\Delta_{0}(Q)\approx 6.43+0.15\,\log Q$, indicating logarithmic corrections to the $1/Q$ law. The ratio $\Delta_{1}/\Delta_{0}$ converges to $\approx 2.37$ as $Q$ grows (computed from the eigenvalue data in Table 4 and confirmed at larger $Q$ values in Figure 2a), confirming that the leading eigenvalue separates from the rest of the spectrum.
@@ 4.1 Outer solution and Fermi sea
Since (20) and the solution varies slowly on the $\mathcal{O}(1)$ scale of the kernel, the convolution integral evaluates to $\approx 2\pi\tilde{\rho}$, and the equation is trivially satisfied for any constant $\tilde{\rho}$.
@@
Numerically, sampling $\tilde{\rho}$ in the outer region ($|\xi|\in[5,Q-5]$) and computing $D(Q)/(2Q)$, both converge to $1/2$ as $Q\to\infty$.
@@ 4.2 Duality with the Love integral equation
The integral equation (27) is closely related to the Love integral equation [64] for the electrostatic potential of two coaxial circular conducting discs separated by a distance $2/Q$. The Love equation reads
@@
$(I-\mathcal{L}_{Q})\tilde{\rho}=\frac{g}{2\pi},\qquad g(\xi)=2/(1+\xi^{2})\,$, (54)
while the Love equation is $(I-\mathcal{L}_{Q})f=1$. The two equations share the same operator $I-\mathcal{L}_{Q}$ but have different right-hand sides: the Lorentzian $g/(2\pi)$ in the lattice NLS case, versus the constant 1 in the Love case.
@@
$D(Q)=\langle\mathbf{1},\tilde{\rho}\rangle=\left\langle(I-\mathcal{L}_{Q})^{-1}\mathbf{1},\,g/(2\pi)\right\rangle=\left\langle f,\,\tfrac{g}{2\pi}\right\rangle=\frac{1}{\pi}\int_{-Q}^{Q}\frac{f(\xi)}{1+\xi^{2}}\,d\xi\,$, (55)